how many milligrams are in 4.6 teragrams? (metric to metric)
1
Expert's answer
2015-11-24T14:56:48-0500
Answer:
1 teragram = 1012 grams
1 gram = 103 milligrams
Therefore, 1 teragram = 1012 * 103 = 1015 milligrams
So, 4.6 teragrams = 4.6 * 1015 milligrams